OpenStudy (anonymous):

Help will Medal! Rita earns $17.00 per hour. If she gets a 6% raise, what will be her new hourly wage?

OpenStudy (anonymous):

not a answer choice

TheSmartOne (thesmartone):

If we have 17$ initially, and then we get a 6& increase that would be: \(\sf \large17+\left(\frac{6}{100}\times 17\right)=?\) Or that cold be written as \(\sf\large \frac{106}{100}\times 17=?\)

OpenStudy (anonymous):

thank you :)

TheSmartOne (thesmartone):

Np :)